OpenStudy (anonymous):

Homework Kim works on Social Studies homework for 2 2/5 h. Then she works on Math homework for 11/4 h. How many hours total does Kim spend doing homework?

OpenStudy (anonymous):



OpenStudy (alexandervonhumboldt2):

2 2/5+1 1/4

OpenStudy (alexandervonhumboldt2):

2+1=3 2/5+1/4=8/20+5/20=13/20 3 13/20

OpenStudy (alexandervonhumboldt2):

is it 1 1/4 or 11/4? @rsdod6

OpenStudy (anonymous):

1 1/4

OpenStudy (alexandervonhumboldt2):

then 3 13/20

OpenStudy (anonymous):

3 13/20 my answer

OpenStudy (alexandervonhumboldt2):

yes